OpenStudy (anonymous):

6.Factorise the following expressions completely. pls check whether my answer is right. @ganeshie8

OpenStudy (anonymous):

a)=-8c^2+39c+5 =8c^2-39c-5 =(c-5)(8c+1)

OpenStudy (anonymous):

but the answer in my book said it's (c-5)(-8c-1)

ganeshie8 (ganeshie8):

-8c^2+39c+5 (40)(-1) = -40 40 - 1 = 39 so you can split the middle term as 40 - 1

ganeshie8 (ganeshie8):

-8c^2+ `39c` +5 -8c^2+ `40c - 1c` +5

ganeshie8 (ganeshie8):

factor the GCF from first two terms and last two terms

ganeshie8 (ganeshie8):

-8c^2+ `39c` +5 -8c^2+ `40c - 1c` +5 -8c(c - 5) - 1(c - 5)

OpenStudy (anonymous):

=(c-5)(-8c-1)

ganeshie8 (ganeshie8):

Yes!

OpenStudy (anonymous):

\(\bf\huge\color{#ff0000}{T}\color{#ff2000}{h}\color{#ff4000}{a}\color{#ff5f00}{n}\color{#ff7f00}{k}~\color{#ffaa00}{Y}\color{#ffd400}{o}\color{#bfff00}{u}\color{#4600ff}{!}\color{#6800ff}{!}\color{#8b00ff}{!}\) @ganeshie8